OpenStudy (anonymous):

an initail investment of $350 is worth $429.20 after six years of continuous compounding. Find the annual interest rate

OpenStudy (anonymous):

continuous interest formula is Pe^(rt) just remember "pert" P stand for principle r is interest rate t is time\[429.20=350e^{(6r)}\]solve for r. You are going to have take the natural log of each side

OpenStudy (anonymous):

thats what confuses me

OpenStudy (anonymous):

???

OpenStudy (anonymous):

ok I'll show you the steps

OpenStudy (anonymous):

Lets get e by itself first\[\frac{429.20}{350}=e^{(6r)}\]\[\ln\frac{429.20}{350}=\ln e^{(6r)}\]ln has base e. \[\ln a^x = b \rightarrow x \ln a=b \]\[\ln e = 1 \]\[\ln\frac{429.20}{350}=6r\ln e\]\[\ln\frac{429.20}{350}=6r\]\[{\ln\frac{429.20}{350} \over 6}=r\]You plug it into a calc

OpenStudy (anonymous):

thanks so much!!